OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 5y = –11 10x + 25y = –55 Is it 0,1,2 Or infinitely many I picked 0 please help me

OpenStudy (anonymous):

A system of equations has only zero, 1, or infinitely many solutions. 1. If the equations are parallel, there are no solutions. They are parallel if, in standard form, the constants are different, like: 2x + 3y = 7 and 2x + 3y = 10 2. if the equations are equal, there are infinite solutions, like: 2x + 3y = 7 and 4x + 6y = 14 notice that if you double each term in the first equation, you will get the second equation, exactly . . . it's the same equation. So the answer would be : ( 2 ), instead of infinity or zero.

OpenStudy (anonymous):

Oh okay thanks i thought it was 0

OpenStudy (anonymous):

You welcome.